# COTR: Compact Occupancy TRansformer for Vision-based 3D Occupancy Prediction

### 04 Dec 2023 v1

@@ Abstract
The autonomous driving community has shown significant interest in 3D occupancy prediction, driven by its exceptional geometric perception and general object recognition capabilities. To achieve this, current works try to construct a Tri-Perspective View (TPV) or Occupancy (OCC) representation extending from the Bird-Eye-View perception. However, compressed views like TPV representation lose 3D geometry information while raw and sparse OCC representation requires heavy but reducant computational costs. To address the above limitations, we propose Compact Occupancy TRansformer (COTR), with a geometry-aware occupancy encoder and a semantic-aware group decoder to reconstruct a compact 3D OCC representation. The occupancy encoder first generates a compact geometrical OCC feature through efficient explicit-implicit view transformation. Then, the occupancy decoder further enhances the semantic discriminability of the compact OCC representation by a coarse-to-fine semantic grouping strategy. Empirical experiments show that there are evident performance gains across multiple baselines, e.g., COTR outperforms baselines with a relative improvement of 8%-15%, demonstrating the superiority of our method.
footnotetext: ${}^{{\dagger}}$ Corresponding author.
@@ 1 Introduction
3D vision perception is now transitioning from Bird-Eye-View (BEV) perception [18, 25, 9, 34, 21] to Occupancy (OCC) perception [10, 36, 40, 24, 33]. The BEV perception excels in 3D object detection tasks due to their unified representation abilities for multi-camera inputs, where the obstruction problem is extremely alleviated in the BEV plane. However, its deficiency in collapsing the height dimensions poses a challenge in preserving the requisite geometric information for a holistic understanding of the 3D scene. To alleviate this issue, [10] proposes a Tri-Perspective View (TPV) representation to delineate 3D scenes. Unfortunately, this introduces a new problem in that the compression along the horizontal dimension leads to significant object overlap.
@@
  • Our method has been embedded into several prevailing backbones, and experiments on the Occ3D-nuScenes and SemanticKITTI datasets show that our approach achieves state-of-the-art performance. What’s more, our method outperforms the backbones with a relative improvement of 8%-15%, as illustrated in Fig. 1.
@@ 3.4 Semantic-aware Group Decoder
Proxy Experiment. To demonstrate that the occupancy feature lacks semantic discriminability, we replace the semantic prediction in our occupancy prediction with the ground-truth label in the corresponding position. As shown in Fig. 3 (b), the mIoU scores are greatly improved, especially for the tail class. This drives us to look for a new approach to greatly enhance the semantic discrimination of the occupancy feature.
@@ 4.1 Experimental Setup
We also integrate our method into two main-stream occupancy models BEVDet4D [8] and TPVFormer [10] to demonstrate the effectiveness of our approach. Unless otherwise specified, all models are trained for 24 epochs using AdamW optimizer [23], in which gradient clip is exploited with learning rate 2e-4.
@@ 4.2 Comparing with SOTA methods
Occ3D-nuScenes. As shown in Table 1, we report the quantitative comparison of existing state-of-the-art methods for 3D occupancy prediction tasks on Occ3D-nuScenes. We integrate our method into TPVFormer and BEVDet4D, and our approach yields significant performance improvements in both geometric completion and semantic segmentation, surpassing the baseline by 3.8%, 1.2% in IoU and 5.1%, 5.2% in mIoU. Notably, our approach based on BEVDet4D utilizes a smaller backbone (ResNet-50) and smaller image input size ($256\times 704$) to achieve mIoU scores of 44.5%, which outperforms both Voxformer [17] (ResNet-101, $900\times 1600$) and SurroundOcc [36] (InternImage-B) by 3.8%. This demonstrates that our method mines more information with a small number of parameters by designing components specifically for 3D occupancy prediction tasks. Compared to the state-of-the-art FBOcc [19], BEVDet4D with COTR surpasses it by 2.3% even without test-time augmentation. In addition, we also scale up the image backbone to SwinTransformer-B [22] and Image Size to $512\times 1408$. The experimental results demonstrate that our method consistently brings performance improvements, even with larger model sizes.
@@ 4.3 Ablation study
The Effectiveness of Each Component. The results are shown in Table 2, we can observe that all components make their own performance contributions. The baseline achieves 70.36% of IoU and 36.01% of mIoU without long-term temporal information. We first integrated the Geometry-aware Occupancy Encoder (GOE) into the baseline model, which brings 0.53% and 1.98% performance gain in IoU and mIoU. It demonstrates that the GOE helps the network to construct a compact 3D occupancy representation with more geometry information, thus aiding semantic recognition. By converting the 3D occupancy task to mask classification with the help of a Transformer decoder (TD), the network’s semantic segmentation capability has been significantly enhanced. By using both GOE and TD, the network can excel in both geometric completion and semantic segmentation, outperforming the baseline by 1.38% of IoU and 4.21% of mIou. Moreover, the Coarse-to-Fine Semantic Grouping further enhances rare object recognition and achieves 41.05% mIoU scores while retaining essentially the same geometry completion capability.
@@ Appendix B Further Experiments
Per-class comparison with SOTA. We report more quantitative details in Table  4 about our experimental results for better comparison with other competitors. Besides TPVFormer [10] and BEVDet4D [8], we also integrate COTR into OccFormer [40]. COTR yields significant performance improvements in both geometric completion and semantic segmentation, surpassing OccFormer by 1.6% in IoU and 3.8% in mIoU. Notably, a conspicuous amelioration primarily resides within small objects and rare objects, demonstrating that our approach can indeed apprehend finer geometric details, and substantially enhance semantic discernibility.
@@
Ablation study on computational cost. As shown in Table  6, our proposed COTR is an efficient approach in which each component does not add a significant amount of computational costs. It’s worth noting that, since we only used the Coarse-to-Fine Semantic Grouping (CFSG) strategy during training and kept only one group during inference, CFSG doesn’t introduce any extra overhead.

### 11 Apr 2024 v2

@@ Abstract
The autonomous driving community has shown significant interest in 3D occupancy prediction, driven by its exceptional geometric perception and general object recognition capabilities. To achieve this, current works try to construct a Tri-Perspective View (TPV) or Occupancy (OCC) representation extending from the Bird-Eye-View perception. However, compressed views like TPV representation lose 3D geometry information while raw and sparse OCC representation requires heavy but redundant computational costs. To address the above limitations, we propose Compact Occupancy TRansformer (COTR), with a geometry-aware occupancy encoder and a semantic-aware group decoder to reconstruct a compact 3D OCC representation. The occupancy encoder first generates a compact geometrical OCC feature through efficient explicit-implicit view transformation. Then, the occupancy decoder further enhances the semantic discriminability of the compact OCC representation by a coarse-to-fine semantic grouping strategy. Empirical experiments show that there are evident performance gains across multiple baselines, e.g., COTR outperforms baselines with a relative improvement of 8%-15%, demonstrating the superiority of our method. The code is available at https://github.com/NotACracker/COTR.
footnotetext: $*{}^{*}$ Equal contribution, ${}^{{\dagger}}$ Corresponding author.
@@ 1 Introduction
3D vision perception is now transitioning from Bird-Eye-View (BEV) perception [24, 18, 26, 9, 35, 21] to Occupancy (OCC) perception [10, 37, 41, 25, 34]. The BEV perception excels in 3D object detection tasks due to their unified representation abilities for multi-camera inputs, where the obstruction problem is extremely alleviated in the BEV plane. However, its deficiency in collapsing the height dimensions poses a challenge in preserving the requisite geometric information for a holistic understanding of the 3D scene. To alleviate this issue, [10] proposes a Tri-Perspective View (TPV) representation to delineate 3D scenes. Unfortunately, this introduces a new problem in that the compression along the horizontal dimension leads to significant object overlap.
@@
  • Our method has been embedded into several prevailing backbones, and experiments on the Occ3D-nuScenes dataset show that our approach achieves state-of-the-art performance. What’s more, our method outperforms the backbones with a relative improvement of 8%-15%, as illustrated in Fig. 1.
@@ 3.4 Semantic-aware Group Decoder
Proxy Experiment. To demonstrate that the occupancy feature lacks semantic discriminability, we replace the semantic prediction in our occupancy prediction with the ground-truth label in the corresponding position. As shown in Fig. 3 (b), the IoU scores are greatly improved, especially for the tail class. This drives us to look for a new approach to greatly enhance the semantic discrimination of the occupancy feature.
@@ 4 Experiments
SurroundOcc* [37] ICCV’23 ResNet-101 $900\times 1600$ 24 65.5 34.6
SurroundOcc + COTR - ResNet-50 $256\times 704$ 24 71.0 39.3
OccFormer* [41] ICCV’23 ResNet-50 $256\times 704$ 24 70.1 37.4
OccFormer + COTR - ResNet-50 $256\times 704$ 24 71.7 41.2
@@ 4.1 Experimental Setup
We also integrate our method into four main-stream occupancy models BEVDet4D [8], TPVFormer [10], SurroundOcc [37] and OccFormer [41] to demonstrate the effectiveness of our approach. Unless otherwise specified, all models are trained for 24 epochs using AdamW optimizer [23], in which gradient clip is exploited with learning rate 2e-4.
@@ 4.2 Comparing with SOTA methods
Occ3D-nuScenes. As shown in Table 1, we report the quantitative comparison of existing state-of-the-art methods for 3D occupancy prediction tasks on Occ3D-nuScenes. We integrate our method into TPVFormer, SurroundOcc, OccFormer and BEVDet4D, and our approach yields significant performance improvements in both geometric completion and semantic segmentation, surpassing the baseline by 3.8%, 5.5%, 1.6%, 1.2% in IoU and 5.1%, 4.7%, 3.8%, 5.2% in mIoU. Notably, our approach based on BEVDet4D utilizes a smaller backbone (ResNet-50) and smaller image input size ($256\times 704$) to achieve mIoU scores of 44.5%, which outperforms both Voxformer [17] (ResNet-101, $900\times 1600$) and SurroundOcc [37] (InternImage-B) by 3.8%. This demonstrates that our method mines more information with a small number of parameters by designing components specifically for 3D occupancy prediction tasks. Compared to the state-of-the-art FBOcc [19], BEVDet4D with COTR surpasses it by 2.3% even without test-time augmentation. In addition, we also scale up the image backbone to SwinTransformer-B [22] and Image Size to $512\times 1408$. The experimental results demonstrate that our method consistently brings performance improvements, even with larger model sizes.
@@ 4.3 Ablation study
The Effectiveness of Each Component. The results are shown in Table 2, we can observe that all components make their own performance contributions. The baseline achieves 70.36% of IoU and 36.01% of mIoU without long-term temporal information. We first integrated the Geometry-aware Occupancy Encoder (GOE) into the baseline model, which brings 0.53% and 1.98% performance gain in IoU and mIoU. By converting the 3D occupancy task to mask classification with the help of a Transformer decoder (TD), the network’s semantic segmentation capability has been significantly enhanced. By using both GOE and TD, the network can excel in both geometric completion and semantic segmentation, outperforming the baseline by 1.38% of IoU and 4.21% of mIou. Moreover, the Coarse-to-Fine Semantic Grouping further enhances rare object recognition and achieves 41.05% mIoU scores while retaining essentially the same geometry completion capability.
@@ 5 Conclusion
Acknowledgements This work is supported by the National Natural Science Foundation of China No.62302167, U23A20343, 62222602, 62206293, 62176224, 62106075, and 62006139, Shanghai Sailing Program (23YF1410500), Natural Science Foundation of Shanghai (23ZR1420400), Science and Technology Commission of Shanghai No.21511100700, Natural Science Foundation of Chongqing, China (CSTB2023NSCQ-JQX0007, CSTB2023NSCQ-MSX0137), CCF-Tencent Rhino-Bird Young Faculty Open Research Fund (RAGR20230121) and CAAI-Huawei MindSpore Open Fund.
@@ Appendix B Further Experiments
Per-class comparison with SOTA. We report more quantitative details in Table  4 about our experimental results for better comparison with other competitors. Besides TPVFormer [10] and BEVDet4D [8], we also integrate COTR into SurroundOcc [37] and OccFormer [41]. COTR yields significant performance improvements in both geometric completion and semantic segmentation, surpassing SurroundOcc and OccFormer by 5.5%, 1.6% in IoU and 4.7%, 3.8% in mIoU. Notably, a conspicuous amelioration primarily resides within small objects and rare objects, demonstrating that our approach can indeed apprehend finer geometric details, and substantially enhance semantic discernibility.
@@
Ablation study on computational cost. As shown in Table  6, our proposed COTR is an efficient approach in which each component does not add a significant amount of computational costs. It’s worth noting that, since we only used the Coarse-to-Fine Semantic Grouping (CFSG) strategy during training and kept only one group during inference, CFSG doesn’t introduce any extra overhead. More metrics are reported in Table 7. All baselines are concurrently equipped with EVT and IVT for a fare comparison.
method Params (M) FLOPs (G) LT Latency (s)
TPVFormer 54.05 972.75 0.59
+ COTR (R50) 53.30 784.85 0.43
BEVDet4D $\dagger$ 35.67 924.07 0.67
+ COTR (R50) 38.87 747.26 0.42
BEVDet4D 35.67 1049.52 1.89
+ COTR (R50) 38.87 747.26 1.43
BEVDet4D 121.28 4195.12 1.67
+ COTR (SwinB) 104.99 3761.06 1.43
Table 7: Ablation study on efficiency. All models are tested on a RTX A6000 GPU. $\dagger$ input size of $256\times 704$, others same as Tab. 1.